OpenStudy (anonymous):

96% of kernals pop when cooked if you cook 125 kernals how many will pop?

OpenStudy (sandra):

in this case you want to multiply the percentage that will be popped by the total number of kernals

OpenStudy (sandra):

so .96 * 125

OpenStudy (sandra):

if you want to do this with fractions, then we know that 96% is 96/100

OpenStudy (sandra):

48/50, and 24/25 (reducing)

OpenStudy (sandra):

so the answer would be 24/25 * 125 = 24*125/25 = 24*25*5/25 = 5*24

OpenStudy (sandra):

so 120 kernels will pop